OpenStudy (anonymous):

A spring has a spring stiffness constant, k, of 440 N/m. How much must this spring be stretched to store 25 J of potential energy?

OpenStudy (anonymous):

potential energy of spring= 1/2 k x^2 25= 1/2 *440* x^2

OpenStudy (anonymous):

0.337099931= x ?

OpenStudy (anonymous):

maybe..

OpenStudy (anonymous):

when I answered your equation...thats tha answer ive got..

OpenStudy (anonymous):

i said maybe...coz that was a lot of calculation there.....and its correct

OpenStudy (anonymous):

okay,,,thank you anyway